# The Jordan canonical form of the Fréchet derivative of a matrix function

### v1

@@ 3.1.2 The Jordan canonical form of $P(J_{m}(\lambda),J_{n}(\mu))$ for $p_{x}\neq 0=p_{y}$
Using Lemma 3.6 (in the scalar case $R=\mathbb{F}$) we see that it suffices to compute the JCF of the nilpotent matrix $N=I_{m}\otimes N_{n}^{r}+N_{m}\otimes I_{n}$. Let $q(x,y)=x+y$; then $N=Q(N_{m},N_{n}^{r})$. Note that $q_{x}=q_{y}=1$ and therefore we can apply the results obtained in the previous section. The JCF of $N_{m}$ is $N_{m}$ itself, while the one of $N_{n}^{r}$ is provided by the following corollary. Corollary 3.7 is a straightforward consequence of [5, Theorem 1.36], or can also be verified by direct inspection:
@@ 3.2 Degenerate case: General bounds
The previous observation is frightening, and we will not delve further into the epic undertaking of a complete solution to the bivariate Jordan problem. Instead, we give some general bounds on the size and number of the Jordan blocks in the degenerate case when (6) fails, based on the the concept of local degree, that we define next. Given a bivariate polynomial $p(x,y)$ as in (2), we define the local degree of $p$ at $(\lambda,\mu)\in\mathbb{F}^{2}$ as the smallest integer $d$ such that there exists a biindex $\alpha=(\beta,\gamma)\in\mathbb{N}^{2}$ with $|\alpha|=\beta+\gamma=d$ and $\hat{\partial}_{z^{\alpha}}p(\lambda,\mu)=\hat{\partial}_{x^{\beta}y^{\gamma}}p(\lambda,\mu)\neq 0$.
@@
When applied to Jordan blocks, [9, Corollary 4.3] is a special case (for $d=1$) of Proposition 3.8.
@@ 4 The Fréchet-Jordan problem
While the bivariate Jordan problem seems generally very hard, the Fréchet-Jordan problem is more structured, and we focus on it in this section with the goal of going beyond the regime of (6), that we already solved more generally. Note that the assumptions of [11, Theorem 2, Theorem 3] bear no relation to the case of a formal Fréchet derivative, and thus our approach will diverge from that of [11].
@@
$H_{k}(x,y):=\sum_{j=0}^{k}x^{j}y^{k-j}=\frac{x^{k+1}-y^{k+1}}{x-y}$ (9)
denote the complete homogeneous symmetric bivariate polynomials of degree $k$. Then, as observed in [9, Theorem 5.1] the case of the formal Fréchet derivative of the polynomial $f(w)$ corresponds to a very special subclass of bivariate polynomials:
@@
Note that (10) is the Bézoutian of $f(w)$ and the constant polynomial $1$. The assumption that $p(x,y)$ has the form (10) imposes a rather rigid structure on the possible values of its formal Hasse derivatives, and thus on the JCF of $P(J_{m}(\lambda),J_{n}(\mu))$; in this section, we exploit this fact to fully solve the Fréchet-Jordan problem. To explain the reason for this helpful rigidity, for a biindex $\alpha=(\beta,\gamma)$ define $z^{\alpha}:=x^{\beta}y^{\gamma}$ and $\hat{\partial}_{z^{\alpha}}:=\hat{\partial}_{x^{\beta}}\hat{\partial}_{y^{\gamma}}$. Proposition 4.1 relates the derivatives of order $\ell-1$ of (10) to the mixed derivatives of order $\ell$, for all $\ell\geq 1$.
@@ 4.1 Distinct eigenvalues
Theorem 4.4 below characterizes the JCF of (4), in the case where $X$ and $Y$ are single Jordan blocks, with distinct eigenvalues.
@@
where $\varphi\in\mathbb{F}[w]$ is the polynomial defined in (11). Moreover, define the integers $a,q,b,r$ as follows. If $k\geq m$ (resp. $h\geq n$), then $q=m$ and $a=0$ (resp. $r=n$ and $b=0$); otherwise, $m=ak+q$ (resp. $n=bh+r$) is the Euclidean division of $m$ by $k$ (resp. of $n$ by $h$). Finally, define the multisets $\{s_{i}\}_{i=1}^{\min(k,m)}$ and $\{t_{j}\}_{j=1}^{\min(h,n)}$ where
@@
Then, for each such pair of Jordan blocks of $X$ and $Y$, and for each of the $kh$ pairs $(i,j)$, with $1\leq i\leq k$ and $1\leq j\leq h$, in the JCF of (4) there are $\min(s_{i},t_{j})$ Jordan blocks of eigenvalue $\frac{f(\lambda)-f(\mu)}{\lambda-\mu}$ and sizes $s_{i}+t_{j}+1-2\min(s_{i},t_{j}),s_{i}+t_{j}+3-2\min(s_{i},t_{j}),\dots,s_{i}+t_{j}-1$.
Proof. By (12) and Lemma 4.3, except possibly for the eigenvalues, the JCF of $P(J_{m}(\lambda),J_{n}(\mu))$ is the same as that of $\varphi(J_{m}(\lambda))\otimes I_{n}-I_{m}\otimes\varphi(J_{n}(\mu))$. The statement then follows by combining Theorem 3.4 and [5, Theorem 1.36].
@@
$J(P)=J_{4}(0)\oplus J_{4}(0)\oplus J_{2}(0)\oplus J_{2}(0).$
@@ 4.2 Equal eigenvalues
It now suffices to observe that, whenever $|\alpha|>0$, it holds $\hat{\partial}_{w^{|\alpha|}}\phi(\lambda)=\hat{\partial}_{w^{1+|\alpha|}}f(\lambda)$.
@@
Let $\lambda\in\mathbb{F}$ and let $1\leq k$ be an integer. If $h(w)\in\mathbb{F}[w]$ satisfies $h(\lambda)\neq 0$, then there exists $r(w)\in\mathbb{F}[[w-\lambda]]$ such that the equation $r(w)]^{k}=h(w)$ holds in $\mathbb{F}[[w]]$.
Proof. Since $h(w)$ is a polynomial, it is clear that it admits an expansion around $\lambda$, that is, $h(w)=\sum_{i=0}^{\deg h}h_{i}(w-\lambda)^{i}$, with $h_{0}\neq 0.$ Let now $\eta\in\mathbb{F}$ be a $k$-th root of $h_{0}$; the existence of $\eta$ is guaranteed because $\mathbb{F}$ is algebraically closed. We then see that it suffices to prove the statement when $h_{0}=1$, for if we can find $\hat{r}(w)$ such that $\hat{r}(w)]^{k}=h(w)/h_{0}$ then we can take $r(w)=\eta\ \hat{r}(w)$ to guarantee $r(w)]^{k}=h(w)$. Note now that the polynomial $y^{k}-1\in\mathbb{F}[x]$ has nonzero discriminant, since it holds (using $\mathrm{char}(\mathbb{F})=0$) $\mathrm{Res}(y^{k}-1,ky^{k-1})=-(-k)^{k}\neq 0$. We can therefore invoke [13, Proposition 6.1.9], from which the statement follows immediately.
We remark that $\mathrm{char}(\mathbb{F})=0$ is crucial for Lemma 4.8, which can fail otherwise. For example, let $\mathrm{char}(\mathbb{F})=p>0$ and take $k=p$, $\lambda=0\in\mathbb{F}$, and $h(w)=1+w\in\mathbb{F}[w]$. Yet, for every $r(w)\in\mathbb{F}[[w]]$,
@@
Suppose that $p(x,y)$ has the form (10) and that it has local degree $d\geq 1$. Then, $P(J_{m}(\lambda),J_{n}(\lambda))-p(\lambda,\lambda)I_{mn}$ is similar to $H_{d}(N_{m},N_{n})$, where $H_{d}(x,y)$ is as in (9).
@@
By Lemma 4.7, we can then write $\phi(w)-\phi(\lambda)=(w-\lambda)^{1+d}h(w)$ for some polynomial $h\in\mathbb{F}[w]$ such that $h(\lambda)\neq 0$. Using Lemma 4.8, take $\hat{r}(w)\in\mathbb{F}[[w-\lambda]]$ to be a formal power series such that $\hat{r}(w)]^{1+d}=h(w)$. Then, defining
@@
$\phi(w)-\phi(\lambda)=[r(w)]^{1+d}\quad\mathrm{and}\quad\hat{\partial}_{w}r(\lambda)=\sqrt[1+d]{h(\lambda)}\neq 0.$
Observe also that there exists a polynomial $s(w)$ such that $s(\lambda)=0$, $\hat{\partial}_{w}s(\lambda)\neq 0$, $s(J_{m}(\lambda))=s(J_{m}(\lambda))$ and $s(J_{n}(\lambda))=r(J_{n}(\lambda))$; indeed, it suffices to take $s(w)$ to be the truncation of the series of $r(w)$ in powers of $(w-\lambda)$, up to the term of order $\max(m,n)-1$. Finally, note the identity
$\frac{\phi(x)-\phi(y)}{x-y}=\frac{r(x)-r(y)}{x-y}\ \frac{[r(x)]^{1+d}-[r(y)]^{1+d}}{r(x)-r(y)}.$ (15)
@@
$P(J_{m}(\lambda),J_{n}(\lambda))-p(\lambda,\lambda)I_{mn}=Q(J_{m}(\lambda),J_{n}(\lambda))\cdot H_{d}(s(J_{m}(\lambda)),s(J_{n}(\lambda)).$
By construction, the only eigenvalue of $Q(J_{m}(\lambda),J_{n}(\lambda))$ is $\hat{\partial}_{w}s(\lambda)\neq 0$. Hence, by Lemma 4.3, $P(J_{m}(\lambda),J_{n}(\lambda))-p(\lambda,\lambda)I_{mn}$ is similar to $H_{d}(s(J_{m}(\lambda)),s(J_{n}(\lambda))$. In turn, since $\hat{\partial}_{w}s(\lambda)\neq 0$ and $s(\lambda)=0$, for $k\in\{m,n\}$ $s(J_{k}(\lambda))$ is similar to $J_{k}(s(\lambda))=N_{k}$ by [5, Theorem 1.36]. At this point, we get the statement by the same argument as in the beginning of Section 2.
Lemma 4.10
Fix $1\leq a,b,d\in\mathbb{N}$ and $c\in\mathbb{N}$. Let $R\in\mathbb{F}^{a\times b}$ be the banded Toeplitz matrix defined as
$R_{ij}=\begin{cases}1\ &\mathrm{if}\ -c\leq j-i\leq d-c;\\ 0\ &\mathrm{otherwise}.\end{cases}$
Suppose that $a\leq b+c\leq a+d$ and that $c\leq d$, and define
$\rho_{1}(a,b,d):=\min(b-1,\max(0,d-a+1))$
and
$\rho_{0}(a,b,c,d):=\begin{cases}\min(b,d+1)\ &\mathrm{if}\ 2\leq a-c\leq d+2;\\ b\ &\mathrm{otherwise}.\end{cases}$
Then, we have
$\mathrm{rank}(R)=\rho(a,b,c,d):=\rho_{0}(a,b,c,d)-\rho_{1}(a,b,d).$
Proof. Denote by $\mathbf{1}_{k}\in\mathbb{F}^{k}$ (resp. $\mathbf{0}_{k}\in\mathbb{F}^{k}$) the all-ones vector (resp. the all-zeros vector). Note that we have $R_{11}=1$ because by assumption $c\leq 0\leq d-c$, and we have $R_{ab}=1$ because by assumption $c\leq b-a\leq d-c$. We now will look for columns equal to $\mathbf{1}_{b}$, since to discount linearly dependent columns we must keep at most one copy of $\mathbf{1}_{b}$. Moreover, we can only keep one column from each pair whose sum is $\mathbf{1}_{b}$.
The condition for the $j$-th column of $R$ to be equal to $\mathbf{1}_{b}$ is $a-c\leq j\leq d-c+1$, and thus there are $\min(b,\max(0,d-a+2))$ such columns in $R$. Hence, $\min(b-1,\max(0,d-a+1))$ of these columns do not contribute to the rank. Moreover, the leftmost column of $R$ is $\begin{bmatrix}\mathbf{1}_{\min(a,c+1)}\\ \mathbf{0}_{\max(0,a-c-1)}\end{bmatrix}$. Hence, no further exclusions are needed if $a\leq c+1$. Otherwise, we note that the vector $\begin{bmatrix}\mathbf{0}_{c+1)}\\ \mathbf{1}_{a-c-1}\end{bmatrix}$ is the $(d+2)$-th column of $R$ provided that $d+2\leq b$ and $d+2\geq a-c$. In this case, we must also exclude this column and any possible further columns to the right. The statement then follows.
@@
Let $H_{d}(x,y)$ be as in (9), assume $m\leq n$, and consider $Z_{m-1}:=H_{d}(N_{m},N_{n})$ If $d\geq m+n-1$, then $Z_{m-1}=0$. Otherwise,
$\mathrm{rank}(Z_{m-1})=\mathfrak{R}(d):=\sum_{k=1+d}^{m+n-1}\rho(u_{k-d},u_{k},c_{k},d)$
where
$c_{k}=\begin{cases}0\ &\mathrm{if}\ k\leq n;\\ k-n\ &\mathrm{if}\ n\leq k\leq n+d;\\ d\ &\mathrm{if}\ k\geq n+d.\end{cases}$
and $\rho$ is the function defined in Lemma 4.10.
Proof. It suffices to deal with the case d<m+n1d<m+n-1. Using the same notation as in the proof of the upper bound in Proposition 3.11, in this case the permuted matrix $T$ has precisely one nonzero block diagonal, and thus $\mathrm{rank}(Z_{m-1})=\sum_{k=d+1}^{m+n-1}\mathrm{rank}(R_{k})$.
By direct inspection, it is clear that each $R_{k}$ is a banded Toeplitz matrix in the form of Lemma 4.10, with $a(k)=u_{k-d}$, $b(k)=u_{k}$, $c(k)=c_{k}$ (if the first column and row in $R_{k}$ correspond to $e_{i_{c}}\otimes f_{j_{c}}$ and $e_{i_{r}}\otimes f_{j_{r}}$, $c_{k}$ describes the offset $i_{c}-i_{r}$), and $d(k)\equiv d$. Manifestly, $0\leq c_{k}\leq d$ for all $k$, and we claim that $a(k)\leq b(k)+c_{k}\leq a(k)+d$ for all $k$, implying the result by Lemma 4.10.
It remains to prove the claim:
  • Let first $k\leq n$. Then $u_{k-d}\leq u_{k}$ because $m\leq n$ (see also the proof of the lower bound in Proposition 3.11). By (7) it is clear that $|u_{k+1}-u_{k}|\leq 1$, implying $u_{k}\leq u_{k-d}+d.$
  • Let $n\leq k\leq n+d$. Then, using (7), $u_{k}=n+m-k$. Moreover, since $k-d\leq n$, either $u_{k-d}=m\geq m-d$ or, if kd<mk-d<m, $u_{k-d}=k-d\geq m-d$. Hence,
    $u_{k-d}\leq m=u_{k}+c_{k}\leq u_{k-d}+d.$
  • Let $k\geq n+d.$ Since $k\geq k+d\geq n$, by (7) we have $u_{k}=n+m-k$ and $u_{k-d}=n+m-k+d$, implying
    $u_{k-d}=n+m-k+d=u_{k}+d\leq u_{k}+2d=u_{k-d}+d.$
@@
Suppose that $p(x,y)$ has the form (10), let $f\in\mathbb{F}[w]$ be the polynomial appearing in (10) and consider the matrix (4). Let $d\geq 1$ be the multiplicity of $\lambda\in\mathbb{F}$ as a root of the polynomial $\phi\in\mathbb{F}[w]$ as defined in Lemma 4.7. For every $k\in\mathbb{N}$, define
$\mathfrak{N}_{k}=\begin{cases}0\ &\mathrm{if}\ k=0;\\ nm-\mathfrak{R}(kd)\ &\mathrm{if}\ 0<kd<m+n-1;\\ mn&\mathrm{otherwise},\end{cases}$
where $\mathfrak{R}(\cdot)$ is the quantity defined in Proposition 4.11. Assume with no loss of generality $m\leq n$. Then, for each pair of Jordan blocks in the JCFs of $X$ and $Y$, having sizes $m$ and $n$ respectively and sharing the same eigenvalue $\lambda$, in the JCF of (4):

### v2

@@ 3.1.2 The Jordan canonical form of $P(J_{m}(\lambda),J_{n}(\mu))$ for $p_{x}\neq 0=p_{y}$
Using Lemma 3.6 (in the scalar case $R=\mathbb{F}$) we see that it suffices to compute the JCF of the nilpotent matrix $N=I_{m}\otimes N_{n}^{r}+N_{m}\otimes I_{n}$. Let $q(x,y)=x+y$; then $N=Q(N_{m},N_{n}^{r})$. Note that $q_{x}=q_{y}=1$ and therefore we can apply the results obtained in the previous section. The JCF of $N_{m}$ is $N_{m}$ itself, while the one of $N_{n}^{r}$ is provided by the following corollary. Corollary 3.7 is a straightforward consequence of [6, Theorem 1.36], or can also be verified by direct inspection:
@@ 3.2 Degenerate case: General bounds
The previous observation is frightening, and we will not delve further into the epic undertaking of a complete solution to the bivariate Jordan problem. Instead, we give some general bounds on the size and number of the Jordan blocks in the degenerate case when (6) fails, based on the the concept of local degree, that we define next. Given a bivariate polynomial $p(x,y)$ as in (2), we define the local degree of $p$ at $(\lambda,\mu)\in\mathbb{F}^{2}$ as the smallest positive integer $d\geq 1$ such that there exists a biindex $\alpha=(\beta,\gamma)\in\mathbb{N}^{2}$ with $|\alpha|=\beta+\gamma=d$ and $\hat{\partial}_{z^{\alpha}}p(\lambda,\mu)=\hat{\partial}_{x^{\beta}y^{\gamma}}p(\lambda,\mu)\neq 0$.
@@
When applied to Jordan blocks, [10, Corollary 4.3] is a special case (for $d=1$) of Proposition 3.8.
@@ 4 The Fréchet-Jordan problem
While the bivariate Jordan problem seems generally very hard, the Fréchet-Jordan problem is more structured, and we focus on it in this section with the goal of going beyond the regime of (6), that we already solved more generally.
@@
$h_{d}(x,y):=\sum_{j=0}^{d}x^{j}y^{d-j}=\frac{x^{d+1}-y^{d+1}}{x-y}$ (9)
denote the complete homogeneous symmetric bivariate polynomials of degree $k$. Then, as observed in [10, Theorem 5.1] the case of the formal Fréchet derivative of the polynomial $f(w)$ corresponds to a very special subclass of bivariate polynomials:
@@
Note that (10) is the Bézoutian of $f(w)$ and the constant polynomial $1$. The assumption that $p(x,y)$ has the form (10) imposes a rather rigid structure on the possible values of its formal Hasse derivatives, and thus on the JCF of $P(J_{m}(\lambda),J_{n}(\mu))$; in this section, we exploit this fact to solve the Fréchet-Jordan problem. To explain the reason for this helpful rigidity, for a biindex $\alpha=(\beta,\gamma)$ define $z^{\alpha}:=x^{\beta}y^{\gamma}$ and $\hat{\partial}_{z^{\alpha}}:=\hat{\partial}_{x^{\beta}}\hat{\partial}_{y^{\gamma}}$. Proposition 4.1 relates the derivatives of order $\ell-1$ of (10) to the mixed derivatives of order $\ell$, for all $\ell\geq 1$.
@@ 4.1 Distinct eigenvalues
Theorem 4.4 below characterizes the JCF of (4), in the case where $X$ and $Y$ are single Jordan blocks, with distinct eigenvalues. In its statement, we exclude the trivial case where $\varphi$ is the zero polynomial, i.e., $f$ is linear and $p$ is constant.
@@
where $\varphi\in\mathbb{F}[w]$ is the polynomial defined in (11) and we assume $\varphi(w)\not\equiv 0$. Moreover, define the integers $a,q,b,r$ as follows. If $k\geq m$ (resp. $h\geq n$), then $q=m$ and $a=0$ (resp. $r=n$ and $b=0$); otherwise, $m=ak+q$ (resp. $n=bh+r$) is the Euclidean division of $m$ by $k$ (resp. of $n$ by $h$). Finally, define the multisets $\{s_{i}\}_{i=1}^{\min(k,m)}$ and $\{t_{j}\}_{j=1}^{\min(h,n)}$ where
@@
Then, for each such pair of Jordan blocks of $X$ and $Y$, and for each of the $\min(k,m)\min(h,n)$ pairs $(i,j)$, with $1\leq i\leq k$ and $1\leq j\leq h$, in the JCF of (4) there are $\min(s_{i},t_{j})$ Jordan blocks of eigenvalue $\frac{f(\lambda)-f(\mu)}{\lambda-\mu}$ and sizes $s_{i}+t_{j}+1-2\min(s_{i},t_{j}),s_{i}+t_{j}+3-2\min(s_{i},t_{j}),\dots,s_{i}+t_{j}-1$.
Proof. By (12) and Lemma 4.3, except possibly for the eigenvalues, the JCF of $P(J_{m}(\lambda),J_{n}(\mu))$ is the same as that of $\varphi(J_{m}(\lambda))\otimes I_{n}-I_{m}\otimes\varphi(J_{n}(\mu))$. The statement then follows by combining Theorem 3.4 and [6, Theorem 1.36].
@@
$J_{4}(0)\oplus J_{4}(0)\oplus J_{2}(0)\oplus J_{2}(0).$
@@ 4.2 Equal eigenvalues
Note that $1+|\alpha|\neq 0$ because $\mathrm{char}(\mathbb{F})=0$. It now suffices to observe that, whenever $|\alpha|>0$, it holds $\hat{\partial}_{w^{|\alpha|}}\hat{\partial}_{w}\phi(\lambda)=\hat{\partial}_{w^{1+|\alpha|}}f(\lambda)$.
The proof of Lemma 4.7 also shows that, for Fréchet derivatives and when $\lambda=\mu$, $p_{x}=p_{y}$ in the notation of Theorem 3.1.
@@
Let $\lambda\in\mathbb{F}$ and let $1\leq k$ be an integer. If $g(w)\in\mathbb{F}[w]$ satisfies $g(\lambda)\neq 0$, then there exists $r(w)\in\mathbb{F}[[w-\lambda]]$ such that the equation $r(w)]^{k}=g(w)$ holds in $\mathbb{F}[[w]]$.
Proof. Since $g(w)$ is a polynomial, it is clear that it admits an expansion around $\lambda$, that is, $g(w)=\sum_{i=0}^{\deg h}g_{i}(w-\lambda)^{i}$, with $g_{0}\neq 0.$ Let now $\eta\in\mathbb{F}$ be a $k$-th root of $g_{0}$; the existence of $\eta$ is guaranteed because $\mathbb{F}$ is algebraically closed. We then see that it suffices to prove the statement when $g_{0}=1$, for if we can find $\hat{r}(w)$ such that $\hat{r}(w)]^{k}=g(w)/g_{0}$ then we can take $r(w)=\eta\ \hat{r}(w)$ to guarantee $r(w)]^{k}=h(w)$. Note now that the polynomial $y^{k}-1\in\mathbb{F}[x]$ has nonzero discriminant, since it holds (using $\mathrm{char}(\mathbb{F})=0$) $\mathrm{Res}(y^{k}-1,ky^{k-1})=-(-k)^{k}\neq 0$. We can therefore invoke [14, Proposition 6.1.9], from which the statement follows immediately.
We remark that $\mathrm{char}(\mathbb{F})=0$ is crucial for Lemma 4.8, which can fail otherwise. For example, let $\mathrm{char}(\mathbb{F})=p>0$ and take $k=p$, $\lambda=0\in\mathbb{F}$, and $g(w)=1+w\in\mathbb{F}[w]$. Yet, for every $r(w)\in\mathbb{F}[[w]]$,
@@
Suppose that $p(x,y)$ has the form (10) and that it has local degree $d\geq 1$. Then, for all $\ell\in\mathbb{N}$, $P(J_{m}(\lambda),J_{n}(\lambda))-p(\lambda,\lambda)I_{mn}]^{\ell}$ is similar to $H_{d}(N_{m},N_{n})^{\ell}$, where $h_{d}(x,y)$ is as in (9).
@@
By Lemma 4.7, we can then write $\phi(w)-\phi(\lambda)=(w-\lambda)^{1+d}g(w)$ for some polynomial $g\in\mathbb{F}[w]$ such that $g(\lambda)\neq 0$. Using Lemma 4.8, take $\hat{r}(w)\in\mathbb{F}[[w-\lambda]]$ to be a formal power series such that $\hat{r}(w)]^{1+d}=g(w)$. Then, defining
@@
$\phi(w)-\phi(\lambda)=[r(w)]^{1+d}\quad\mathrm{and}\quad\hat{\partial}_{w}r(\lambda)=\sqrt[1+d]{g(\lambda)}\neq 0.$
In addition, define $s(w)$ to be the polynomial obtained by truncating the series $r(w)$ in powers of $w-\lambda$ up to the term of order m+n1m+n-1. This implies, in particular, that $s(\lambda)=0$, $\hat{\partial}_{w}s(\lambda)\neq 0$, $s(J_{m}(\lambda))=r(J_{m}(\lambda))$, and $s(J_{n}(\lambda))=r(J_{n}(\lambda))$. Finally, note the identity
$\left[\frac{\phi(x)-\phi(y)}{x-y}\right]^{\ell}=\left[\frac{r(x)-r(y)}{x-y}\right]^{\ell}\ \left[\frac{[r(x)]^{1+d}-[r(y)]^{1+d}}{r(x)-r(y)}\right]^{\ell}.$} (15)
@@
$\left[P(J_{m}(\lambda),J_{n}(\lambda))-p(\lambda,\lambda)I_{mn}\right]^{\ell}=\left[Q(J_{m}(\lambda),J_{n}(\lambda))\right]^{\ell}\cdot\left[H_{d}(s(J_{m}(\lambda)),s(J_{n}(\lambda))\right]^{\ell}.$}
By construction, the only eigenvalue of $Q(J_{m}(\lambda),J_{n}(\lambda))]^{\ell}$ is $(\hat{\partial}_{w}s(\lambda))^{\ell}\neq 0$. Hence, by Lemma 4.3, $P(J_{m}(\lambda),J_{n}(\lambda))-p(\lambda,\lambda)I_{mn}]^{\ell}$ is similar to $H_{d}(s(J_{m}(\lambda)),s(J_{n}(\lambda))]^{\ell}$. In turn, since $\hat{\partial}_{w}s(\lambda)\neq 0$ and $s(\lambda)=0$, for $k\in\{m,n\}$ $s(J_{k}(\lambda))$ is similar to $J_{k}(s(\lambda))=N_{k}$ by [6, Theorem 1.36]. At this point, we get the statement by the same argument as in the beginning of Section 2.
By Theorem 4.9, the problem of computing the JCF of (4) when $X,Y$ are two Jordan blocks of size $m$ and $n$ and with the same eigenvalues is therefore reduced to computing the ranks of $H_{d}(N_{m},N_{n})^{\ell},\ell\in\mathbb{N}$, where $d$ is the local degree of (10). Indeed, it is a classical result [15] that, for a nilpotent matrix $Z$, the number of Jordan blocks of size $s$ and eigenvalue 0 is given by
$2\dim\ker Z^{s}-\dim\ker Z^{s-1}-\dim\ker Z^{s+1}.$
If $d\ell\geq m+n-1$, then clearly $H_{d}(N_{m},N_{n})^{\ell}=0$. Otherwise, using the same notation as in the proof of the upper bound in Proposition 3.11, we can see that the nilpotent matrix $H_{d}(N_{m},N_{n})^{\ell}$ is permutation similar to a matrix $T$ that, when $d\ell<m+n-1$, has precisely one nonzero block diagonal. The block elements of this nonzero block diagonal are the banded Toeplitz rectangular matrices $R_{k}\in\mathbb{F}^{(u_{k}-\ell d)\times u_{k}}$, for $d\ell+1\leq k\leq m+n-1$, and where $u_{k}$ are the integers defined by (7). Specifically, $R_{k}$ represents the action of $H_{d}^{\ell}(N_{m},N_{n})$ when restricted to a linear map $V_{k}/V_{k-1}\rightarrow V_{k-\ell d}/V_{k-\ell d-1}$. In particular, the nonzero entries of $R_{k}$ corresponds to the coefficients of the homogeneous polynomial $h_{d}(x,y)^{\ell}=\left(\sum_{i=0}^{d}x^{i}y^{d-i}\right)^{\ell}$. Hence, $R_{k}$ depends only on the quintuple $(m,n,d,\ell,k)$, and in particular, $R_{k}\in\mathbb{Z}^{(u_{k}-\ell d)\times u_{k}}$ is an integer banded Toeplitz matrix. (Note that, since $\mathbb{F}$ has characteristic zero, then necessarily $\mathbb{Z}\subsetneq\mathbb{F}$.) By some tedious, but straightforward, computations, one can see that this description is equivalent to Definition 4.10 below.
Definition 4.10
Fix the positive integers $m,n,d,\ell,k$, such that $m\leq n$, and $d\ell+1\leq k\leq m+n-1$. Let the integers $u_{k}$ be defined as in (7). Furthermore, for $0\leq i\leq\ell d$, define $\gamma_{i}$ by
$\sum_{i=0}^{\ell d}\gamma_{i}z^{i}=(1+z+\dots+z^{d})^{\ell},$
and set $\gamma_{i}=0$ for every other value of $i$.
Then, we define the integer Toeplitz matrices $R_{k}\in\mathbb{Z}^{u_{k-\ell d}\times u_{k}}$ by setting $(R_{k})_{ij}=\gamma_{j-i+c_{k}}$ where
$c_{k}=\begin{cases}0\ &\mathrm{if}\ k\leq n;\\ k-n\ &\mathrm{if}\ n\leq k\leq n+\ell d;\\ \ell d\ &\mathrm{if}\ k\geq n+\ell d.\end{cases}$
If we denote the first column and row in $R_{k}$, respectively, by $e_{i_{c}}\otimes f_{j_{c}}$ and $e_{i_{r}}\otimes f_{j_{r}}$, then in Definition 4.10 the quantity $c_{k}=i_{c}-i_{r}$ describes the offset in the banded matrix $R_{k}$. In Proposition. 4.11 we annotate some futher properties of $R_{k}$.
@@
Fix $m,n,d,\ell,k$ and, without loss of generality, assume $m\leq n$. Then, the banded Toeplitz matrix $R_{k}$ in Definition 4.10 satisfies the following properties:
  1. 1.
    $0\leq c_{k}\leq\ell d$;
  2. 2.
    $u_{k-\ell d}\leq u_{k}+c_{k}\leq u_{k-\ell d}+\ell d$;
  3. 3.
    The top left element of $R_{k}$ is equal to $\gamma_{c_{k}}>0$;
  4. 4.
    The bottom right element of $R_{k}$ is equal to $\gamma_{u_{k}-u_{k-\ell d}+c_{k}}>0$;
  5. 5.
    Let $F_{s}\in\mathbb{F}^{s\times s}$ be the flip matrix (the matrix whose elements are $1$ on the main antidiagonal and 0 elsewhere). Then, $F_{u_{k-\ell d}}R_{k}F_{u_{k}}=R^{T}_{\ell d+m+n-k}$, for all $d\ell+1\leq k\leq m+n-1$. In particular, $R_{k}$ and $R_{\ell d+m+n-k}$ have the same rank.
Proof.
  1. 1.
    Immediate by the definition of $c_{k}$.
  2. 2.
    We distinguish three cases.
    • Let first $k\leq n$. Then $u_{k-\ell d}\leq u_{k}$ because $m\leq n$ (see also the proof of the lower bound in Proposition 3.11). By (7) it is clear that $|u_{k+1}-u_{k}|\leq 1$, implying $u_{k}\leq u_{k-\ell d}+\ell d.$
    • Let $n\leq k\leq n+\ell d$. Then, using (7), $u_{k}=n+m-k$. Moreover, since $k-\ell d\leq n$, either $u_{k-\ell d}=m\geq m-\ell d$ or, if $k-\ell d<m$, $u_{k-\ell d}=k-\ell d\geq m-\ell d$. Hence,
      $u_{k-\ell d}\leq m=u_{k}+c_{k}\leq u_{k-\ell d}+\ell d.$
    • Let $k\geq n+\ell d.$ Since $k\geq k-\ell d\geq n$, by (7) we have $u_{k}=n+m-k$ and $u_{k-\ell d}=n+m-k+\ell d$, implying
      $u_{k-\ell d}=n+m-k+\ell d=u_{k}+\ell d\leq u_{k}+2\ell d=u_{k-\ell d}+\ell d.$
  3. 3.
    By the first item, $0\leq c_{k}\leq\ell d$ and thus $\gamma_{c_{k}}$ is a coefficient of the polynomial $(1+z+\dots+z^{d})^{\ell}$. By the multinomial theorem, it follows that $\gamma_{c_{k}}$ is the sum of some positive integers.
  4. 4.
    By the second item $0\leq u_{k}-u_{k-\ell d}+c_{k}\leq\ell d$. We can then follow the same argument as in the previous item.
  5. 5.
    For simplicity of notation, set $h:=\ell d+m+n-k$ and $k^{\prime}=k-\ell d$. It follows from (7) that the sizes of $R_{k}$ and $R_{h}$ are compatible with the statement. We leave as an exercise to the reader to verify, from (7) and Definition 4.10, that $c_{h}+u_{k^{\prime}}=c_{k}+u_{k}$ holds for all $k$ in the range of definition of $R_{k}$. Hence,
    $(F_{u_{k^{\prime}}}R_{k}F_{u_{k}})_{ij}=(R_{k})_{u_{k^{\prime}}+1-i,u_{k}+1-j}=\gamma_{i-j+u_{k}-u_{k^{\prime}}+c_{k}}=\gamma_{i-j+c_{h}}=(R_{h})_{ji}.$
@@
Suppose that $p(x,y)$ has the form (10), let $f\in\mathbb{F}[w]$ be the polynomial appearing in (10) and consider the matrix (4) assuming without loss of generality that $m\leq n$. Let $d\geq 1$ be the multiplicity of $\lambda\in\mathbb{F}$ as a root of the polynomial $\phi\in\mathbb{F}[w]$ as defined in Lemma 4.7. For every positive integer $\ell$ and for every $d\ell+1\leq k\leq m+n-1$, let $\rho(k,m,n,d,\ell)$ be the rank of the Toeplitz matrix $R_{k}$ of Definition 4.10. Let
$\mathfrak{N}_{s}:=\begin{cases}0\ &\mathrm{if}\ s=0;\\ nm-\sum_{k=ds+1}^{n+m-1}\rho(k,m,n,d,s)\ &\mathrm{if}\ 0<sd<m+n-1;\\ mn&\mathrm{otherwise}.\end{cases}$
Then, for each pair of Jordan blocks in the JCFs of $X$ and $Y$, having sizes $m$ and $n$ respectively and sharing the same eigenvalue $\lambda$, in the JCF of (4):